OpenStudy (anonymous):

Initially, a 1.50-liter compressible container of 0.35 moles of a gas exerts a pressure of 740 millimeters of mercury at a temperature of 275 kelvin. What is the pressure when the container is compressed to 0.50 liters, the moles of gas reduces to 0.20 moles, and the temperature changes to 32 degrees Celsius? 1.1 x 103 mm Hg 1.4 x 103 mm Hg 1.5 x 102 mm Hg 3.9 x 102 mm Hg

OpenStudy (anonymous):

@Somy @Abhisar @Abmon98

OpenStudy (anonymous):

@Abmon98 do you get it or no

OpenStudy (abmon98):

Use PV=nRT 740(mm Hg)*1.50(L)=0.35(moles)*R*275(K) R=740(mm Hg)*1.50(L)/0.35(moles)*275(K)=11.53 mm Hg/mol-1 K Ideal gas constant is known. use PV=nRT again Change degrees celsium to Kelvin by adding to 273. P*0.50=0.20*11.53*(273+32) P=(.20*11.53*(273+32)/0.50=1406.66 mm Hg

OpenStudy (anonymous):

so it would be B right @Abmon98

OpenStudy (abmon98):

yes

OpenStudy (anonymous):

thank you :) @Abmon98

OpenStudy (abmon98):

your welcome :D